Assume that $200 is deposited today, two years from now, four years from now, six years from now, and eight years from now in a savings account which earns 10% annual interest, what is the balance at the end of year 9?

1517

708

1100

1670
Answer #1

Solution: 1670
Working:

Year   Deposit   Interest @ 10%   Value     
1   200   20   220     
2   220   22   242     
3   242+200 = 442   44   486     
4   486   49   535     
5   535+200 = 735   73   808     
6   808   81   889     
7   1089   109   1198     
8   1198   120   1318     
9   1318+200 = 1518   152   1670